a)If the acetate to acetic acid ratio of a solution is five, what is its pH? The pKa of acetic acid is 4.75.

b)What is the dominant ionic form of phosphoric acid in a solution at pH 8.2?
Answer #1

pH of Buffer solution is calculated by using Henderson's equation as

pH = pKa + log [ salt] / [ Acid]

pH = pKa + log [Acetate]  / [Acetic acid]

pH = 4.75 + log 5

pH = 4.75 + 0.6990

pH = 5.45

ANSWER : pH = 5.45

Consider dissociation of Phosphoric acid in water.

H3PO4 (aq) + H2O (l) H3O + (aq) + H2PO4- (aq)

For above reaction, K1 = [H3O +] [H2PO4-] / [H3PO4 ]

We have, pK1 = - log K1 = 2.148

H2PO4- (aq) + H2O (l) H3O + (aq) + HPO42- (aq)

For above reaction, K2 = [H3O +] [HPO42-] / [H2PO4- ]

We have, pK2 = - log K 2 = 7.198

HPO42- (aq) +  H2O (l) H3O + (aq) + PO43- (aq)

For above reaction, K2 = [H3O +] [PO43-] / [HPO42-]

We have, pK3 = - log K 3 = 12.375

We know that, 1) if pH < pKa then acidic form predominates and if pH > pKa basic form predominates.

pH of solution is 8.2 . It is in between pKa values 7.198 and 12.375. Hence, predominant form of phosphoric acid is HPO42-.